In a recent wind storm, Hobbit Company's computer system was knocked out and they lost a portion of their accounting records. The IT manager at Hobbit was able to retrieve the following information:

Sales $1,500,000

Net Operating Income $850,000

Total Assets

Return on Investment 0.15

Residual Income

What is the residual income for the Hobbit Company if the firms cost of capital is 13%?

Question 8 options:

$113,333

$127,501

$195,027

$650,000
Answer #1

Solution:
Given in the question
Sales = $1500000
Net operating income = $850000
Return on investment = 0.15
First we will calculate Average operating asset = Net operating income/Return on investment = ($850000/0.15) = 5666667
Now residual income can be calculated as
Residual income = Net operating income - (Cost of capital*Average operating asset) = 850000 - 0.13*5666667 = 850000 - 736666.7 = $113333.3 or 113333
So its correct answer is A i.e. residual income = $113333